OpenStudy (anonymous):

Write an equation of a line that passes through (9,-5) and (3,-5)

OpenStudy (anonymous):

An equation of a line is y=mx + c wherein m=slope or gradient and c is the y-intercept. So technically, you have to achieve this format. The easiest way is by following the "point-slope" formula which is: \[y - y_1 = m(x - x_1)\] The points are: (9, -5) and (3, -5) Wherein y1 = -5, y2 = -5, x1 = 3, x2 = 9 Substitute. Next, find the slope, using the formula: slope = change in y/ change in x Substitute and solve.

OpenStudy (anonymous):

thnk u